# Multi-Modal Hypergraph Enhanced LLM Learning for Recommendation

### v1

@@ 3.6. Objective Optimization
where and are the balance factors for and , respectively for the two hypergraphs. The last term is the regularization with coefficient to prevent overfitting. represents the Bayesian Personalized Ranking (BPR) loss (Rendle et al., 2012), which employs a pairwise approach to optimize the prediction of higher scores for observed items compared to unobserved ones, defined as:
@@
4. EXPERIMENTS
4.1. Experimental Settings
4.1.1. Datasets.
To evaluate our proposed method, we conduct experiments on three public benchmark datasets from Amazon 222Datasets are available at http://jmcauley.ucsd.edu/data/amazon/links.html(McAuley et al., 2015): Toys and Games (Toys), Beauty, and Sports and Outdoors (Sports), with dataset statistics in Table 2. Each dataset includes visual and textual modalities. For the textual modality, we concatenate titles, brands, categories, and descriptions, processing them through BERT 333https://huggingface.co/google-bert/bert-base-uncased (Devlin et al., 2018) to obtain 768-dimensional features. For the visual modality, images are sourced from the original URLs, with 512-dimensional features extracted via VIT444https://huggingface.co/openai/clip-vit-base-patch32 (Radford et al., 2021). Following prior work (Kang and McAuley, 2018), we split each user sequence by using the last interaction as the test set, the penultimate as the validation set, and all preceding interactions as the training set.
@@ 4.1. Experimental Settings
4.1.3. Baseline methods.
@@
To validate our model’s effectiveness, we compare it with various state-of-the-art recommender models, including i) the general collaborative filtering methods LightGCN (He et al., 2020), ii) multimodal GNN-based CF methods (e.g., VBPR (He and McAuley, 2016), GRCN (Wei et al., 2020), LATTICE (Zhang et al., 2021), SLMRec (Tao et al., 2022), FREEDOM (Zhou and Shen, 2023), BM3 (Zhou et al., 2023)), and iii) LLM-based methods Collm (Zhang et al., 2023b) and LLaRA (Liao et al., 2024). For a fair comparison, we use the same pre-trained LLM backbone (LLaMA-2 7B(Touvron et al., 2023)) for all LLM-based models, and the same textual prompts used in Fig.2. Specifically, Collm and LLaRA utilize LightGCN (He et al., 2020) and SASRec (Kang and McAuley, 2018) as their pre-trained recommender embeddings, respectively. The results are summarized in Table 1.
4.1.4. Implementation details.
@@
We implemented all the experiments using PyTorch on an RTX 4090 GPU with 24GB memory. For a fair comparison with previous works (Zhou and Shen, 2023; Zhou et al., 2023), the recommender dimension is set to 64 for all pre-trained models. Our model is trained using the Adam optimizer (Kingma and Ba, 2014), with learning rates and batch sizes set at 0.0001 and 1024 for pre-training, and 0.0003 and 7 for fine-tuning, respectively. We use the Xavier initializer (Glorot and Bengio, 2010) for initializing the user and item embeddings. The hyperparameters and are optimized by searching from 0.1 to 0.9 in 0.2 increments. The number of k-nearest neighbors varies from 1 to 30 in increments of 5. The contrastive learning temperature is adjusted from 0.1 to 1 in 0.1 increments. More details are presented in Sec. 4.4.
4.2. Performance Comparison
The experimental results, presented in Table 1, reveal the following key observations:
@@ 4.2. Performance Comparison
0.0542 | 0.0221 | 0.0820 | 0.0342 | 0.0786 | 0.0319 | |
HeLLM | 0.0617 | 0.0279 | 0.1021 | 0.0465 | 0.1027 | 0.0460 |
4.3. Ablation Study
To systematically evaluate the contribution of each component in HeLLM, we conduct a comprehensive ablation study, as summarized in Table 3. Our analysis encompasses both the recommendation pre-training phase and the LLM fine-tuning stage HeLLM, where key architectural elements are selectively removed to assess their individual impact.
Effectiveness of and Multi-View Hypergraph Learning. We first conducted experiments for variants of the recommender pre-training by removing different hypergraph modules and contrastive learning. The omission of the U2U hypergraph (w/o-U2U) leads to the most significant performance drop, particularly in the Sports and Toys datasets. This confirms the pivotal role of the U2U hypergraph in modeling complex shared preferences among users. Similarly, excluding I2I hypergraph (w/o-I2I) results in noticeable degradation, especially in the Beauty dataset, underscoring the importance of item-item multimodal semantic relations. The impact of contrastive learning (w/o-CL) is also evident, with a decline across all domains, reaffirming its role in enforcing robust representation learning. Notably, our HeLLMpre consistently outperforms all GNN-based baselines in the pre-training phase, verifying its effectiveness in capturing multi-view graph topological information. These findings highlight the advantage of hypergraphs for capturing shared user preferences and intricate multimodal semantic correlation among items.
@@ 4.3. Ablation Study
4.4. Impact Study of Hyperparameters
@@
4.4.1. Impact of Synergistic Contrastive Learning.
As shown in Fig. 4, and denote the significance of the U2U and I2I hypergraphs, respectively. The results show that when both and are set to zero, disabling contrastive learning, the performance noticeably degrades. However, with both and active (e.g., at values like and ), the model demonstrates substantial performance gains. This improvement highlights the essential role of synergistic contrastive learning in the pre-training phase, which effectively enhances feature distinguishability.
4.4.2. Impact of Top-.
@@
4.4.3. Impact of Contrastive Learning (CL) Temperature Varied .
@@
4.4.4. Impact of Regularization Coefficient .
@@
4.4.5. Impact of Embedding Dimensionality .
As shown in Fig. 3(j)-(l), performance improves with increasing embedding dimensions up to 256, after which it plateaus or slightly declines. This trend indicates that while larger dimensions enhance representational capacity, excessive embedding size may introduce redundancy, highlighting the importance of selecting an optimal dimension to balance informativeness and complexity.
5. Conclusion
In this study, we identify two key objectives for advancing LLM-based recommender systems: (1) enhancing the expressive power of pre-trained representations, and (2) effectively integrating rich graph structure information to strengthen the LLM’s ability to perceive relational patterns. To this end, we propose HeLLM, a framework that integrates multi-view hypergraph signals and sequential patterns into LLMs. We construct user and item hypergraphs to uncover shared interests and capture complex higher-order multimodal semantic correlations. Experiments on multiple benchmarks validate the effectiveness of our proposed method.

### v2

@@ III-F Objective Optimization
where and are the balance factors for and , respectively for the two hypergraphs. The last term is the regularization with coefficient to prevent overfitting. represents the Bayesian Personalized Ranking (BPR) loss [75], which employs a pairwise approach to optimize the prediction of higher scores for observed items compared to unobserved ones, defined as:
@@
IV Theoretical Analysis of Complexity
To elucidate the computational efficiency of our proposed hypergraph-based framework, we provide a detailed theoretical analysis encompassing both the hypergraph construction and convolutional propagation stages. For clarity and generality, we take the item-to-item (I2I) hypergraph as a representative case, as it captures the typical complexity patterns of other hypergraph variants (e.g., user-to-user). This analysis quantifies both time and space complexities, offering insights into the scalability and practicality of the proposed design.
IV-A Hypergraph Construction Complexity
The total construction time consists of three stages: multimodal pairwise similarity computation , top-K neighbor selection , and normalization over non-zero entries . Thus, the total time complexity is:
The construction of the initial similarity matrix incurs a space complexity of due to its dense nature. However, following -nearest neighbor (KNN) pruning, the matrix is sparsified, reducing the space requirement to , which significantly improves memory efficiency.
IV-B Hypergraph Convolution Complexity
To enable efficient computation, we precompute the sparse matrix , where each row has at most non-zero entries due to the KNN pruning, resulting in a practical storage size of . This design transforms the hypergraph convolution into a single sparse-dense matrix multiplication , with time complexity and space complexity . This matches the complexity of standard GCN or HGNN layers while reducing redundant computation via precomputation.
IV-C User-to-User (U2U) Hypergraph Complexity
The complexity of U2U hypergraph construction and convolution is analogous to I2I hypergraph, with construction time complexity replace . The resulting hypergraph matrix remains sparse and supports a similar convolution process. In this setting, the sparsity reflects the average number of user–item interactions per user. For simplicity, we omit the constant scale factors such as the number of modalities and layers, which would linearly scale the total complexity. The computational complexity and experimental details for hypergraph construction and convolution are provided in Sec. V-D.
V EXPERIMENTS
V-A Experimental Settings
V-A1 Datasets
To evaluate our proposed method, we conduct experiments on three public benchmark datasets from Amazon 222Datasets are available at http://jmcauley.ucsd.edu/data/amazon/links.html[76]: Sports (Sports and Outdoors), Beauty, and Toys (Toys and Games) with dataset statistics in Table I. Each dataset includes visual and textual modalities. For the textual modality, we concatenate titles, brands, categories, and descriptions, processing them through BERT 333https://huggingface.co/google-bert/bert-base-uncased [77] to obtain 768-dimensional features. For the visual modality, images are sourced from the original URLs, with 512-dimensional features extracted via VIT444https://huggingface.co/openai/clip-vit-base-patch32 [78]. Following prior work [18], we split each user sequence by using the last interaction as the test set, the penultimate as the validation set, and all preceding interactions as the training set.
@@ V-A2 Metrics
V-A3 Baseline methods
| Baseline | Sports | Beauty | Toys | |||||||||
| Conventional Recommendation Methods | ||||||||||||
| LightGCN [2] | 0.0232 | 0.0120 | 0.0411 | 0.0165 | 0.0426 | 0.0208 | 0.0669 | 0.0269 | 0.0273 | 0.0138 | 0.0409 | 0.0172 |
| VBPR [30] | 0.0298 | 0.0150 | 0.0485 | 0.0197 | 0.0436 | 0.0216 | 0.0724 | 0.0289 | 0.0376 | 0.0196 | 0.0565 | 0.0244 |
| GRCN [32] | 0.0303 | 0.0152 | 0.0502 | 0.0202 | 0.0454 | 0.0225 | 0.0758 | 0.0305 | 0.0411 | 0.0214 | 0.0610 | 0.0264 |
| LATTICE [33] | 0.0310 | 0.0155 | 0.0506 | 0.0204 | 0.0468 | 0.0234 | 0.0760 | 0.0307 | 0.0459 | 0.0228 | 0.0683 | 0.0284 |
| SLMRec [41] | 0.0312 | 0.0156 | 0.0507 | 0.0205 | 0.0458 | 0.0227 | 0.0754 | 0.0301 | 0.0436 | 0.0223 | 0.0675 | 0.0284 |
| BM3 [40] | 0.0316 | 0.0160 | 0.0517 | 0.0210 | 0.0446 | 0.0217 | 0.0722 | 0.0286 | 0.0402 | 0.0204 | 0.0618 | 0.0258 |
| FREEDOM [36] | 0.0314 | 0.0157 | 0.0507 | 0.0205 | 0.0516 | 0.0248 | 0.0805 | 0.0321 | 0.0467 | 0.0232 | 0.0700 | 0.0291 |
| LGMRec [79] | 0.0321 | 0.0164 | 0.0519 | 0.0213 | 0.0495 | 0.0240 | 0.0795 | 0.0316 | 0.0451 | 0.0228 | 0.0715 | 0.0294 |
| PGL [45] | 0.0302 | 0.0152 | 0.0486 | 0.0198 | 0.0449 | 0.0220 | 0.0719 | 0.0288 | 0.0436 | 0.0212 | 0.0651 | 0.0266 |
| Large Language Models-based Methods | ||||||||||||
| Collm [10] | 0.0349 | 0.0182 | 0.0525 | 0.0226 | 0.0681 | 0.0393 | 0.0960 | 0.0463 | 0.0652 | 0.0376 | 0.0905 | 0.0439 |
| LLaRA [11] | 0.0367 | 0.0203 | 0.0562 | 0.0252 | 0.0667 | 0.0380 | 0.0925 | 0.0445 | 0.0683 | 0.0371 | 0.0944 | 0.0437 |
| (Ours) | 0.0329 | 0.0168 | 0.0542 | 0.0221 | 0.0527 | 0.0269 | 0.0820 | 0.0342 | 0.0498 | 0.0246 | 0.0786 | 0.0319 |
| HeLLM (Ours) | 0.0412 | 0.0227 | 0.0617 | 0.0279 | 0.0696 | 0.0384 | 0.1021 | 0.0465 | 0.0751 | 0.0404 | 0.1081 | 0.0488 |
@@ V-A3 Baseline methods
To validate our model’s effectiveness, we compare it with various state-of-the-art recommender models, including i) the general collaborative filtering methods LightGCN [2], ii) multimodal GNN-based CF methods (e.g., VBPR [30], GRCN [32], LATTICE [33], SLMRec [41], BM3 [40], FREEDOM [36], LGMRec [79], PGL [45]) and iii) LLM-based methods Collm [10] and LLaRA [11]. The detailed descriptions of each method are provided below.
LightGCN [2] removes feature transforms and nonlinearity from GCNs, relying solely on linear neighborhood aggregation and layer combination for recommendation.
VBPR [30] firstly integrates visual features into a scalable factorization model to uncover users’ visual preferences. This pioneering work established the foundation for multimodal recommendation.
GRCN [32] adaptively prunes false-positive edges in interaction graphs via refining layers, then applies graph convolution.
LATTICE [33] mines latent item-item structures from multimodal content by learning modality-aware item graphs to enhance item representation.
SLMRec [41] introduces self-supervised learning by augmenting multimodal content through feature dropout and masking to enhance item representations.
BM3 [40] bootstraps latent contrastive views via dropout and jointly optimizes interaction reconstruction and cross-/intra-modal alignment to learn representations.
FREEDOM [36] simplifies LATTICE by fixing item-item graphs and denoising the user-item interaction graph via degree-sensitive edge pruning.
LGMRec [79] decouples collaborative and multimodal signals via local graphs and leverages a hypergraph module to model global user interests.
PGL [45] extracts the principal subgraphs with global and local aware operators to capture both co-occurrence patterns and personalized signals.
Collm [10] enhances LLM-based recommendation by treating collaborative signals as an independent modality and aligning them with LLM inputs via a learnable mapping module, while using LoRA for fine-tuning.
LLaRA [11] enhances sequential recommendation by aligning ID-based behavioral embeddings with the LLM input space through hybrid prompting.
To ensure a fair comparison, all LLM-based models adopt the same pre-trained backbone, LLaMA-2 7B [80], and share the same textual prompt templates as illustrated in Fig.2. Specifically, CoLLM and LLaRA incorporate item embeddings derived from LightGCN [2] and SASRec [18], respectively. All LLM methods are optimized using the Cross-Entropy loss. The results are summarized in Table II.
V-A4 Implementation details
@@
We implemented all the experiments using PyTorch on an RTX 4090 GPU with 24GB memory. For a fair comparison with previous works [36, 40], the recommender dimension is set to 64 for all pre-trained models. Our model is trained using the Adam optimizer [81], with learning rates and batch sizes set at 0.0001 and 1024 for pre-training, and 0.0003 and 7 for fine-tuning, respectively. We use the Xavier initializer [73] for initializing the user and item embeddings. The hyperparameters and are optimized by searching from 0.1 to 0.9 in 0.2 increments. The number of k-nearest neighbors varies from 1 to 30 in increments of 5. The contrastive learning temperature is adjusted from 0.1 to 1 in 0.1 increments. More details are presented in Sec. V-E.
V-B Performance Comparison
The experimental results, presented in Table II, reveal the following key observations:
@@ V-B Performance Comparison
TABLE III: Resource Consumption Across All Stages of Our Framework on Three Benchmark Datasets Dataset Users Items Sparsity U-I Time U-I Mem U2U Time U2U Mem I2I Time I2I Mem Pretrain Time Pretrain Mem FT Trainable FT Time FT Mem Sports 35,598 18,357 99.96% 281.6s 8.59MB 53.6s 195.69MB 54.2s 22.92MB 2h57min 16,151MB 1.66% 15h33min 22,227MB Beauty 22,363 12,101 99.94% 80.3s 5.87MB 14.0s 101.67MB 33.7s 139.53MB 41min 7,515MB 1.66% 9h49min 22,537MB Toys 19,412 11,924 99.94% 58.5s 4.91MB 10.2s 57.48MB 33.1s 79.53MB 3h38min 6,633MB 1.66% 8h6min 22,047MB -
@@
| 0.0542 | 0.0221 | 0.0820 | 0.0342 | 0.0786 | 0.0319 | |
| HeLLM | 0.0617 | 0.0279 | 0.1021 | 0.0465 | 0.1081 | 0.0488 |
V-C Ablation Study
To systematically evaluate the contribution of each component in HeLLM, we conduct a comprehensive ablation study, as summarized in Table IV. Our analysis encompasses both the recommendation pre-training phase and the LLM fine-tuning stage HeLLM, where key architectural elements are selectively removed to assess their individual impact. From this, we can observe that:
Effectiveness of and Multi-View Hypergraph Learning. We first conducted experiments for variants of the recommender pre-training by removing different hypergraph modules and contrastive learning. The omission of the U2U hypergraph (w/o-U2U) leads to the most significant performance drop, particularly in the Sports and Toys datasets. This confirms the pivotal role of the U2U hypergraph in modeling complex shared preferences among users. Similarly, excluding I2I hypergraph (w/o-I2I) results in noticeable degradation, especially in the Beauty dataset, underscoring the importance of item-item multimodal semantic relations. The impact of contrastive learning (w/o-CL) is also evident, with a decline across all domains, reaffirming its role in enforcing robust representation learning. Notably, our consistently outperforms all GNN-based baselines in the pre-training phase, verifying its effectiveness in capturing multi-view graph topological information. These findings highlight the advantage of hypergraphs for capturing shared user preferences and intricate multimodal semantic correlation among items.
@@ V-C Ablation Study
V-D Computation Cost and Efficiency
Table III provides a detailed overview of the computational cost and resource consumption of our framework across all pipeline stages. For each dataset, we report the number of users and items, interaction sparsity, and the time and memory required to construct the user-item (U-I) graph as well as the user-user (U2U) and item-item (I2I) hypergraphs. These graph and hypergraph construction processes benefit from our sparsity-aware design, which reduces the storage requirement from an initial dense matrix to after KNN pruning, leading to significantly lower memory overhead (e.g., less than 200MB for U2U construction on the Sports dataset). All hypergraph matrices are precomputed once and reused throughout the training pipeline, effectively minimizing redundant computation and memory overhead.
In addition, we further measure GPU-side resource requirements during the pretraining (PT) and fine-tuning (FT) phases, including wall-clock time, peak GPU memory usage, and the proportion of trainable parameters. During FT, only 1.66% of parameters are updated with a batch size of 7, demonstrating that the proposed PEFT strategy enables efficient adaptation while keeping the memory footprint manageable. For example, our largest dataset, Sports (35,598 users and 18,357 items), can be processed within 3 hours of pretraining using 16GB GPU memory, and fine-tuning takes approximately 15 hours with 22GB GPU memory on a single NVIDIA RTX 4090 (24GB).
V-E Impact Study of Hyperparameters
@@
V-E1 Impact of Synergistic Contrastive Learning
As shown in Fig. 4, and control the contribution of user-user (U2U) and item-item (I2I) hypergraph-based contrastive signals, respectively. When both coefficients are set to zero, effectively disabling the contrastive learning modules, the model suffers a significant drop in performance, indicating that merely relying on structural encoders without contrastive regularization limits the model’s ability to learn discriminative representations. Conversely, enabling both and yields notable improvements, even with small values such as and , suggesting that introducing lightweight contrastive supervision already yields strong gains. This supports our hypothesis that the complementary signals captured by U2U and I2I views play a synergistic role in guiding pretraining. Moreover, the performance surface exhibits a relatively smooth basin centered around low to moderate and , implying robustness to hyperparameter perturbation. However, excessively large values (e.g., and ) do not yield further gains, and may even lead to slight degradation, likely due to over-regularization overriding structural cues.
V-E2 Impact of Top-
@@
V-E3 Impact of Contrastive Learning (CL) Temperature Varied
@@
V-E4 Impact of Regularization Coefficient
@@
V-E5 Impact of Embedding Dimensionality
As shown in Fig. 3(j)-(l), increasing the embedding dimensionality initially improves performance, with optimal results typically observed at d=256, after which it plateaus or slightly declines. This trend indicates that while higher dimensions enhance representational capacity, overly large embeddings introduce redundancy and unnecessary computational overhead. Thus, selecting an appropriate dimensionality is essential for balancing informativeness and efficiency.
VI Conclusion
In this study, we identify two key objectives for advancing LLM-based recommender systems: (1) enhancing the expressive power of pre-trained representations, and (2) effectively integrating rich graph structure information to strengthen the LLM’s ability to perceive relational patterns. To this end, we propose HeLLM, a novel framework that enhances LLMs’ graph reasoning capabilities for multimodal recommendation. During pre-training, user and item hypergraphs uncover shared interests and refine multimodal correlations, while a synergistic contrastive learning strategy enhances feature discriminability. In the fine-tuning phase, HeLLM employs mapped feature embeddings and graph pooling to inject enriched graph-structured and multimodal information into the LLM, strengthening its reasoning ability for recommendation tasks. Additionally, sequence-based embeddings capture temporal behavioral patterns, further enriching the model’s representation capacity. Experiments on three benchmarks validate the effectiveness of our proposed method.